OpenStudy (anonymous):

A body was found at 6.am outdoors on a day when the temperature was 50 degrees. The medical examiner found the temperature of the body to be 66 degrees. What was the approximate time of death? Use newtons law of cooling, with k= 0.1947 T(t) = Ta + (To - Ta)e^kt

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

I'm a little tired and in need of sleep here, sorry i can't do a step by step guide, but i would try and find out rate of change of temp per hr from 66 deg to 50 degs, and take it from the 6am.

OpenStudy (anonymous):

Thanks soo much

OpenStudy (anonymous):

np,hope I helped

OpenStudy (anonymous):

Would it be 5 am, 2 am, 3 am, or midnight

OpenStudy (anonymous):

i still need help

OpenStudy (anonymous):

hmm give me a few mins see if i can work it out for you, very tired here sorry

OpenStudy (anonymous):

okay thanks sorry but really need help

OpenStudy (anonymous):

can you tell mw which variables the t0,ta are please ? i'm guessing Ta = time aprox

OpenStudy (anonymous):

to=intial temp ta=ambient temp

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

tt = 50 + (16)e(^-0.1947 x change in time) => tt -50 /16 = e(^-0.1947 x change in time) => change in time = ln (tt - 50/16) / -0.1947 (k) hmm without a change in time, or a temp of body at time of death I don't see how to do this, very sorry :C I did find a useful tool that may help though http://www.endmemo.com/physics/coollaw.php

OpenStudy (anonymous):

but if i just cheat here :} by putting in 0.5 hrs as time diff i got 69 degs as final, only 3 of the 66 hmm..

OpenStudy (anonymous):

so

OpenStudy (anonymous):

time diff of 0.2 hrs gets you 66.6 final temp btw

OpenStudy (anonymous):

5 am?

OpenStudy (anonymous):

yeah I think so :)

OpenStudy (anonymous):

damnit sorry i know I know can be more helpful with sleep

OpenStudy (anonymous):

it was midnight

OpenStudy (anonymous):

that was the anser

OpenStudy (anonymous):

if you sub in 6 hrs time diff, the final temp doesn't give me 66 degs :L

OpenStudy (anonymous):

idk thats what it said

OpenStudy (anonymous):

ahh , forgot i swapped the sign :D

OpenStudy (anonymous):

yeah now if i sub in -6 :P not 6 you get final temp of, you guess it: 66 degs,s o midnight is right :}

OpenStudy (anonymous):

sorry lol